An inverted regular square pyramid has an altitude of 12m and with top
edge of 3m is fully–filled with water.
1. What is the volume of the water in m3?
1
Expert's answer
2020-11-03T13:56:11-0500

"V = 1/3*S*h = 1/3 * 3^2 *12 =\\= 3*12 = 36 cm^3"